If 1/7*(1/a-1/b) = (a-b)*(1/-5)
where a>b>0. What is the value of (a)(b)?
Please help

Answer by Boreal(15235)

1/7*(1/a-1/b) = (a-b)*(1/-5)
(b-a)/ab=(a-b)(-7/5), multiplying through and putting the fraction over a common denominator.
Now multiply both sides by -1.
(b-a)(-1)=(a-b), which you want
(a-b)/ab=(a-b)(7/5)
cancel the (a-b)
(1/ab)=(7/5)
ab=5/7 by inverting both sides